Ribbon A is 1/3m long. It is 2/5m shorter than ribbon B. What's the total length of the two ribbons?
Answer by Theo(13342)

ribbon A is 1/3 meters long.
it is 2/3 meters shorter than ribbon B.
what's the total length of the two ribbons.

the length of ribbon A is 1/3 meters long.
the length of ribbon B is 1/3 meters long + 2/3 meters long = 3/3 meters long = 1 meters long.
the total length of both ribbons is 1 meters long plus 1/3 meters long = 1 and 1/3 meters long.